Listed below are the lead concentrations in μ​g/g measured in different traditional medicines. Use a

0.01 significance level to test the claim that the mean lead concentration for all such medicines is less than 13 μ​g/g.

6     

8     

3.5     

10.5     

3    

21.5     

4    

3.5    

18     

8

Determine the test statistic
Answer #1

Ans:

From above data:

n=10

sample mean=8.6

sample standard deviation=6.415

Test statistic:

t=(8.6-13)/(6.415/sqrt(10))

t=-2.169

df=10-1=9

p-value=tdist(2.169,9,1)=0.0291

As,p-value is greater than 0.01,so we fail to reject null hypothesis.

There is not sufficient evidence to support the claim that the mean lead concentration for all such medicines is less than 13 μ​g/g.